OpenStudy (anonymous):

what is the sum of the geometric sequence -4, 24, -144 if there are 8 terms

OpenStudy (anonymous):

@timo86m @DAMONSANCHEZ @dinisha

OpenStudy (anonymous):

first we find the common ratio by dividing the 2nd term by the first term. 24 / -4 = -6 so our common ratio is -6 now we use the sum formula : sn = a1 (1 - r^n) / (1 - r) n = sum of the terms we want to find = 8 a1 = first term = -4 r = common ratio = -6 now we sub s8 = -4 (1 - (-6^8) / (1 - (-6) s8 = -4 (1 + 1679616) / (1 + 6) s8 = -4 (1679616) / 7 s8 = - 6718464/7 s8 = -959781

OpenStudy (anonymous):

life saver thank you! can u help with another please please

OpenStudy (anonymous):

I can try

OpenStudy (anonymous):

what are the explicit equation and domain for a geometric sequence with a first term of 4 and a second term of -8

OpenStudy (anonymous):

explicit equation : an = a1 (r ^ n -1) a1 = 4 r = -8/4 = -2 now sub an = 4 (-2^n-1) I do not know hw to do the domain

OpenStudy (anonymous):

ok thanks.. maybe u can do antoher one